The number of gram molecules of chlorine in 6.022×10^25 hydrogen chloride molecule is

Grade:10

2 Answers

Jogender
7 years ago
Given= 6.022×10^25To find = gram molecules which means molesSo simply moles=no of molecules given÷avogadro noSo 6.022×10^25÷6.022×10^23We will get answers as 100 moles or gram molecules
Pooja
3 years ago
 
Given= 6.022×10^25To find = gram molecules which means molesSo simply moles=no of molecules given÷avogadro noSo 6.022×10^25÷6.022×10^23We will get answers as 100 moles or gram atoms. Since Chlorine is diatomic therefore 100/2 = 50 gram molecules